OpenStudy (anonymous):

A triangle has a side with length 6 feet and another side with length 8 feet. The angle between the sides measures 73 degrees. Find the area of the triangle. Round your answer to the nearest tenth. A) 1752.0 ft^2 B) 7.0 ft ^2 C) 45.9 ft^2 D) 23.0 ft^2

jimthompson5910 (jim_thompson5910):

A = b*c*sin(angle) A = 6*8*sin(73) A = 48*sin(73) A = ???

OpenStudy (anonymous):

45.9, so it's C! Thank you!

jimthompson5910 (jim_thompson5910):

you got it